2. 120 points A wagon of mass 48.6 kg has an initial velocity of + 4.76 m/s. If the wagon comes to rest after moving 5.83 m, what was its constant acceleration?
Answer #1

Mass of wagon = m = 48.6 Kg

Initial velocity = u = 4.76m/s

Final Velocity = v = 0 m/s

V2-U2 = 2as

a = 4.762/2s = 4.762/(2 x 5.83) = 1.94m/s2

hence constant acceleration was 1.94 m/s2